AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 01:28:10 2025
Author
*
:
Subject
*
:
<p> </p> <p>Beam energy - 1.361 MeV/u</p> <p>17:12</p> <p>Run 28 Started - beam on target</p> <p>Detectors IN (left arm pot = 36.5 mm, right arm pot = 70.5 mm)</p> <p>Beam position (vertical offset -9mm, horizontal offset -7.5mm)</p> <p>Ions circulate in the ring for 3 seconds before they are dumped. Total time between injections is 12 seconds.</p> <p>Target density apporx 4.2e11 atoms/cm2, pressures in the ring are below 1e-10 mbar. (Attachment 1+2)</p> <p>Leakage current plot (attachment 3) - change in leakage current is consistent with temperature variation from ML online plot.</p> <p>18:18</p> <p>FEE temperatures OK - aida09 is slightly hot - 65.06 C so slightly above vertex temp value of 65C and stays this way over several reloads (attachment 4)</p> <p>System wide checks OK</p> <p>Statistics (attachments 5+6) - aida09 and aida12 slightly high stats but still <1K. Check load does not resolve.</p> <p>Tape server OK - writing ~730 kB/sec</p> <p>Merger OK</p> <p>Monitoring code XY plots and Ey energy histograms (attachments 7+8)</p> <p>18:53 All checks OK</p> <p>21:58</p> <p>All system checks ok. Temperatures and leakage current ok. Merger ok.</p> <p>Attatchment 9: Rates and stats - all below 1k</p> <p>Attatchment 10: Temperatures</p> <p>Attatchment 11-14: Waveforms layouts 1,2,7,8 respectively look okay</p> <p>Attatchment 15: Leakge current</p> <p>Attatchment 16: lowenergyXY plots</p> <p>Vacuum pressure is ok.</p> <p>23:01</p> <p>system wide checks ok. Temperatures ok. Leakage current ok. Merger ok all went green. Waveforms same as 10pm check.</p> <p>Statistics - aida09 and aida12 went over 1k (attach 17), but reloaded and settled down again seconds appart (attatch 18). Reloaded again at 23:14 and stats were aove 1k again (attatch 19).</p> <p>Attatchment 20 is the leakage current, still ok. </p> <p>Vacuum pressure ok (attatchment 21) - rises after small bump down from the regular value.</p> <p>01:00</p> <p>System wide checks ok. Temperatures ok. Leakage current ok (see attatchment 22). Merger ok all went green. Waveforms the same as 10pm check.</p> <p>Statistics - aida09 and aida12 occasionally just over 1k when reloading a few times, but goes below mostly.</p> <p>Vacuum pressure ok, stablised after small bump around 10ish (attatchment 23)</p> <p>03:00</p> <p>Ring vacuum suffered from long-term creep, especially in YR08. Tweaked nozzle position. Pressure improves, target density slightly affected. Attachment 24-25</p> <p>Leakage current and temperature indicate we have an increase in temperature in the ring. This is confirmed by vacplot. Attachment 26-27</p> <p>Merger continues OK 28</p> <p>Rates and Stats show we have no beam. Going to the lab 29-30.</p> <p> </p> <p>R28_22 NO BEAM</p> <p> </p> <p>03:45</p> <p>No beam in the ring, we are missing all the spills. Called HKR.</p> <p>Detectors out in case beam manipulation is required.</p> <p>Stat plot looked OK. 31</p> <p>04:35</p> <p>DAQ continues OK on run R28_24, still no beam</p> <p>Temperatures slightly lower - OK</p> <p>System wide checks - OK</p> <p>Stats - OK</p> <p>Merger - OK</p> <p>05:44</p> <p>Beam is back. It had been away since 1:30 am due to a fault in the new EC power supply. Fault reason unclear.</p> <p>Detectors back in (left 36.54, right 70.49 mm)</p> <p> </p> <p>R28_26 BEAM ON</p> <p>(Beam likely off from R28_ 21-25 inclusive)</p> <p> </p> <p>Temps as before</p> <p>Stats now OK - att 32</p> <p>Merger OK - att 33, 34</p> <p>Stat spectra OK - att 35. Note slope in ASICS (Rutherford?)</p> <p>Leakage current trending up - likely temperature related.</p> <p>Ring pressure OK</p> <p>Gas target 4.14E11</p> <p>Beam at injection around 7.5E6. Could be higher. Asked for optimisation later today assuming current configuration holds.</p> <p>06:59</p> <p>Temperatures OK</p> <p>System checks OK</p> <p>Stats looking good in all FEEs except aida01 and aida15</p> <p>Leakage current has levelled off - looks OK</p> <p>Ring pressure looks OK and is trending down - target density 4.1e11</p> <p>Merger and tape server OK</p> <p>08:30</p> <p>System checks all OK</p> <p>Temperatures OK</p> <p>Merger / tape server OK - data being written R28_29</p> <p>Target density 4.1e11, ring pressure stable</p> <p>09:00</p> <p>Attachment 38 Spectrum browser stats for all FEEs (log scale)</p> <p>Attachment 39-42 Spectrum browser *L histograms for select channels for each detector. MIDAS data is consistent with inline monitoring plots.</p> <p>System checks OK</p> <p>Beam still on - target density and ring pressure remain stable</p> <p>11:28 We have access to the voltage dividers for the electron cooler. The voltage is -763.625 V. This corresponds to an electron energy of 746.795 eV and a beam energy of 1.361 MeV/u using ML ion calc.</p> <p>This is above the resonance energy of 1.3435 MeV/u. We will now end the run and set the cooler voltage to this ion energy.</p> <p>Run ended. Run 29 started while setting beam energy.</p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown